A chemist adds 226 mL of a solution of Na2SO4 to 500 mL of 1.308 M AgCH3CO2. Assume the reaction goes to 100% completion and that Na2SO4 is the limiting reagent. If 76.81 g of Ag2SO4 (mM = 311.802 g/mol) is produced, what was the concentration, in mol/L, of the Na2SO4 solution that was added
1
Expert's answer
2020-09-21T06:34:57-0400

Na2SO4+2AgCH3CO2=Ag2SO4+2CH3COONa

1.in the case condition the "limiting reagent" is Na2SO4. that is, Na2SO4 was completely reacted

2.Based on the above reaction, 1 mole of Na2SO4 produces 1 mole (311.8gr)of Ag2SO4.

3.We find how many moles of Na2SO4 were used to produce 76.81 g of Ag2SO4:

1mole Na2SO4— 311.8 gr Ag2SO4

x mole Na2SO4 — 76.81 gr Ag2SO4

x= 76.81*1/311.8=0.246 mole Na2SO4

4.We find the concentration of Na2SO4.To do this, we divide the amount of Na2SO4 (mol) by the volume (liters):

0.246 mol/ 0.226 litrs = 1.09 mol/L ≈1.1 mol/l

ANSWER: 1.1mol/L